At what angular frequency is the voltage amplitude across (a)resistor, (b) inductor, (c) capacitor, in an LCR series circuit at maximum value?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear ravi

current in LC R circuit

i = V/[R2 + (1/wC -wL)2]1/2

 voltage across resistor  = iR  = VR /[R2 + (1/wC -wL)2]1/2

voltage across inductor  = iwL  = VwL /[R2 + (1/wC -wL)2]1/2

voltage across capacitor  = i/wc  = V /wC[R2 + (1/wC -wL)2]1/2

 

for maximum value differentiate each w.r.t w and put equal to zero

 for resister w=1/√LC

for inductor  =√[2C/(2L3 -CL2R2)]

for capacitor=√[(2L-CR2)/2L2 ]

 

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin